0.08x + 0.07(15,000 - x)=1,110 solve the equation by first clearing it of decimals
Answer by longjonsilver(2297)

0.08x + 0.07(15,000 - x)=1,110
8x + 7(15000-x) = 111000
8x + 105000-7x = 111000
x + 105000 = 111000
x = 6000

jon.